A television repairman finds that the time spent on his jobs has an exponential distribution with a mean of 30 minutes . If he repairs sets in order in which they come in , and if arrival of sets follows a Poisson distribution approximately with an average rate of 10 per 8 hours day , what is the repairman's expected ideal time each day, how many jobs are ahead of the average set just brought in?


1
Expert's answer
2021-04-28T10:11:38-0400

"\lambda=\dfrac{10}{8}=1.25" sets per hour, and "\mu=\dfrac{1}{30}(60)=2" sets per hour


(a) Expected idle time of repairman each day = Number of hours for which the repairman

remains busy in an 8-hour day (traffic intensity) is given by

              

        "8(\dfrac{\lambda}{\mu})=8(\dfrac{1.25}{2})=5 hours"


Hence, the idle time for a repairman in an 8-hour day will be: (8 – 5) = 3 hours.


(b)Expected (or average) number of TV sets in the system


    "E(L_s)=\dfrac{\lambda}{\mu-\lambda}=\dfrac{1.25}{2-1.25}=\dfrac{1.25}{0.75}=1.66 \approx 2( Approx)".